# How to Synthesize Text Data without Model Collapse?

### v1

@@ 7 Conclusion
With the growing prevalence of generative AI models like ChatGPT (Achiam et al., 2023) and Stable Diffusion (Rombach et al., 2021), when training next-generation AI models, it will be inevitable to use a mixture of synthetic data and human-produced data. Therefore, we focus on two key questions: (1) What is the impact of synthetic data on language model pre-training, and what are the underlying causes? (2) How can we prevent model collapse and synthesize high-quality data? We found that synthetic data can impair the effectiveness of pre-training when mixed with human-produced data, leading to non-iterative model collapse. Statistical analysis reveals that synthetic data suffers from significant distribution gaps and overly concentrated n-gram features. To address this, we propose token-level editing instead of relying purely on synthetic data. Specifically, we perform token resampling guided by a trained prior. Theoretically, our method can prevent model collapse. Experimentally, our approach demonstrates improvements over the source data across pre-training, continual pre-training, and supervised fine-tuning.
@@ A.2 Proof of Theorem 2
Appendix B More Results of Human and Synthetic Data Mixture Training
@@
We provide more training results for the human and synthetic data mixture. The main results and analysis can be found in Sec 2.1. Except for GPT-2 pretraining, we also use the OLMo models (Groeneveld et al., 2024) for further experiments.
@@
As shown in Figure 8, the training loss continues to decrease as the amount of synthetic data increases, which is consistent with GPT-2 pretriaing in Figure 2. More synthetic data can lead to better fitting. However, a lower loss does not necessarily mean a better model. As illustrated in Figure 2B and 8, models that fits better perform worse in real world tasks.
@@
Furthermore we follow Maini et al. (2024) to conduct more experiments including PPL results on 22 validation sets of Pile (Gao et al., 2020a) and general understanding tasks. The additional results in Table 1011 and 1 are consistent with our findings. Specifically, the PPL increases as the proportion of purely synthetic data grows, while the performance on downstream tasks similarly exhibits a gradual decline with the increase in synthetic data.
@@
Appendix C Experiment Settings
@@ C.1 Training
We follow Cheng et al. (2024b) to conduct continual pre-training on biomedicine, finance, and math domains. Specifically, PubMed Abstracts from the Pile are utilized as the pre-training corpora for the biomedicine domain. For the finance domain, financial news data covering over 7,000 stocks from May 2022 to May 2023 is collected using the FinGPT framework. We continue pre-training OLMo-1B and LLaMA-3-8B on each domain. For implementation, we utilized the official training framework for OLMo-1B, leveraging Fully Sharded Data Parallel (FSDP) for continual pretraining. For LLaMA, we adopted the LLaMA-Factory framework to carry out the continual pretraining process. Our experiments was primarily conducted on OLMo-1B and Llama-3-8B models, with Llama-3-8B utilizing LoRA (Low-Rank Adaptation) for parameter-efficient fine-tuning. The data and evaluation are given in this repo [^4] https://github.com/microsoft/LMOps/tree/main/adaptllm. We conducted the continual pretraining on a total of 1B tokens.
@@
We used the Llama-Factory (Zheng et al., 2024) framework to fine-tune Llama-3-8B. As for general instruction tuning tasks, we adopt instruction tuning datasets from (Xia et al., 2024)  [^5] https://huggingface.co/datasets/princeton-nlp/less_data, including CoT (Wei et al., 2022) , FLAN v2 (Longpre et al., 2023), and Open Assistant 1 (Kopf et al., 2023). As for code-related reasoning tasks, we utilize OSS-Instruct-75K  [^6] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K and Evol-Instruct-110K  [^7] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K. These datasets provide sufficient diversity for verification on fine-tuning.
C.2 Evaluation
@@
Table 12: PPL results of GPT-2 124M pretraining on mixture of human and synthetic data.
@@
Appendix D Discussion
D.1 What is Coverage Collapse?
‘Coverage collapse’ refers to a phenomenon in which the distribution of synthetic data covers a significantly narrower range of values compared to human data, even when the data sizes are identical. For instance, as shown in Figure 3, the PPL range of synthetic data is limited to $[0,14]$, whereas the PPL range of human data extends from $[0,100]$. Despite this disparity, the overall coverage, represented by the area under the distribution curves, remains the same. This significant distribution gap is what we define as ‘coverage collapse.’
D.2 How Does the DSIR Work?
@@
Table 14: Comparison of different synthetic data methods.
Method Data Type Approach Result
D.3 Non-autoregressive Token Replacement May Compromise Text Coherence.
@@
D.4 Comparison with Pure Synthetic Data and Reformat Methods
D.5 Must We Assume the Data is 100% Human-authored?
@@
Appendix E Potential Applications and Future Work

### v2

@@ 6 Conclusion
With the growing prevalence of generative AI models, when training next-generation AI models, it will be inevitable to use a mixture of synthetic data and human-produced data. Therefore, we focus on two key questions: (1) What is the impact of synthetic data on language model pre-training, and what are the underlying causes? (2) How can we prevent model collapse and synthesize high-quality data? We found that synthetic data can impair the effectiveness of pre-training when mixed with human-produced data, leading to non-iterative model collapse. Statistical analysis reveals that synthetic data suffers from significant distribution gaps and overly concentrated n-gram features. We propose token-level editing instead of relying purely on synthetic data. Specifically, we perform token resampling guided by a trained prior. Theoretically, our method can prevent model collapse. Our approach demonstrates improvements over the source data across pre-training, continual pre-training, and supervised fine-tuning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgement
This work is sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402).
@@ A.2 Proof of Theorem 2
Figure 6: Token distribution across different probability ranges in BioMed dataset.
Figure 7: Pre-training loss of GPT-2 Small (124M) on human (Dolma (Soldaini et al., 2024)) and synthetic (Cosmopedia (Ben Allal et al., 2024)) data. As the proportion of synthetic data increases, the model’s loss decreases.
Appendix B More Related Work
Phi-1/2 (Gunasekar et al., 2023) demonstrate that the synthetic data can boost training efficiency and performance compared to raw data in language model pre-training. Furthermore, Feng et al. (2024) introduce a verifier to filter synthetic samples, theoretically avoiding model collapse. Liu et al. (2024); Tan et al. (2024) highlight that synthetic data will play a crucial role in the development of AI. For example, synthetic data can be used to construct highly specialized datasets, enhancing the performance of downstream tasks. Trinh et al. (2024) utilize synthetic math data to train a 125M language model, which successfully solved 25 out of 30 selected problems from the International Mathematical Olympiad (IMO) problem set. Zhang et al. (2024) develop a biomedical instruction dataset that was used to train specialized bio-models, enabling them to excel in answering questions related to medical exams and clinical scenarios. Eldan & Li (2023) introduce a novel synthetic dataset and evaluation paradigm that enables small language models to generate coherent, diverse, and grammatically sound stories. As outlined above, in the post-training stages of LLMs, synthetic data enhances downstream task performance and aligns foundation models with humans. And Maini et al. (2024) propose rephrasing the pre-training data into a Wikipedia or Q/A style to achieve better alignment with downstream tasks. Synthetic data is a powerful tool for training. Our approach is also based on synthetic data methods. Instead of sampling data solely based on this prior, we modify the data using the prior as a guide.
Bertrand et al. (2023) develop a rigorous framework to demonstrate the importance of real data in maintaining the stability of iterative training. Ferbach et al. (2024) theoretically demonstrate that the impact of data curation can be formalized as an implicit preference optimization mechanism. Kazdan et al. (2025) reveal the detailed training dynamics of model collapse under three different training workflows. Of course, there are also some remarkable studies that successfully used synthetic data. Wang et al. (2022) propose the Self-Instruct data generation framework, enhancing instruction-following capabilities. Ulmer et al. (2024) employ the self-talk method to generate high-quality data. ReST (Gulcehre et al., 2023) uses a policy model to generate datasets and then employs offline RL to fine-tune LLMs on generated datasets. Singh et al. (2023) demonstrate that self-training with binary feedback filtering can reduce reliance on real data. Alemohammad et al. (2023) demonstrate that without enough fresh real images, future generative models will gradually decline. Briesch et al. (2023) illustrates that real data in the iterative training process can slow the decline of LLMs, but cannot fully prevent it. Martínez et al. (2023) shows that the quality and diversity of generated images degrade over time.
Appendix C More Discussion of Main Results
As shown in Table 3, our method shows consistent improvements over the source data across OLMo-1B and LLaMA-3-8B. For instance, in the Biomedicine domain, the average score for OLMo-1B increased from 36.63 to 40.89 with ToEdit, while LLaMA-3-8B saw an increase from 54.13 to 56.48. Table 2 further supports the effectiveness of our approach in pre-training. The average performance of OLMo-1B increases from 32.75 to 33.11, reflecting improved generalization capabilities. While the improvement is modest, the consistent trend across tasks like PIQA, BoolQ, and ARC-c highlights the broader applicability of our method. As for SFT results in Table 4, using both the original and edited data, the results indicate a small but consistent improvement. Specifically, ToEdit improves original FLAN v2, with average performance increasing from 70.18 to 70.65. For Natural Instructions, the average performance of LLaMA-3-8B improves from 69.34 to 69.70, with gains in tasks like Winogrande and SIQA. These improvements demonstrate the adaptability of our method to instruction-tuning tasks. For code-related tasks, the improvements demonstrate better reasoning and code comprehension.
Appendix D Comparison with Pure Synthetic Data and Reformat Methods
Definition and Characteristics of Synthetic Data
Table 10: Comparison of different synthetic data methods.
Method Data Type Approach Result
@@ Appendix D Comparison with Pure Synthetic Data and Reformat Methods
Appendix E More Results of Human and Synthetic Data Mixture Training
@@
We provide more training results for the human and synthetic data mixture. The main results and analysis can be found in Sec 3.1. Except for GPT-2 pre-training, we also use the OLMo models (Groeneveld et al., 2024) for further experiments.
@@
As shown in Figure 9, the training loss continues to decrease as the amount of synthetic data increases, which is consistent with GPT-2 pre-training in Figure 2. More synthetic data can lead to better fitting. However, a lower loss does not necessarily mean a better model. As illustrated in Figure 2B and 9, models that fits better perform worse in real world tasks.
@@
Furthermore we follow (Maini et al., 2024) to conduct more experiments including PPL results on 22 validation sets of Pile (Gao et al., 2020a) and general understanding tasks. The additional results in Table 911 and 1 are consistent with our findings. Specifically, the PPL increases as the proportion of purely synthetic data grows, while the performance on downstream tasks similarly exhibits a gradual decline with the increase in synthetic data.
@@
Appendix F Experiment Settings
@@ F.1 Training
Table 12: Case Study.
Before (source) After (edited) Changes
Construct a function using PHP language that applies lexical analysis on a provided text string to analyze the individual, non-repeated words elements present. Construct a function using PHP language that applies lexical analysis on a provided text string to quantify unique words. “analyze” $\rightarrow$ “quantify”
Test with provided string, \$str = ’Greetings, Planet Earth!’. Test with provided string, \$str = ’Greetings, Planet Earth!’. No changes.
Implements wordCount to remove punctuation, convert text to lowercase, split into words, and count unique words. Implements wordCount to remove punctuation, convert text to lowercase, split into words, and calculate unique words. “count” $\rightarrow$ “calculate”
Returns {’greetings’: 1, ’planet’: 1, ’earth’: 1}. Returns {’greetings’: 1, ’planet’: 1, ’earth’: 1}. No changes.
@@
We follow (Cheng et al., 2024b) to conduct continual pre-training on biomedicine, finance, and math domains. Specifically, PubMed Abstracts from the Pile are utilized as the pre-training corpora for the biomedicine domain. For the finance domain, financial news data covering over 7,000 stocks from May 2022 to May 2023 is collected using the FinGPT framework. We continue pre-training OLMo-1B and LLaMA-3-8B on each domain. For implementation, we utilized the official training framework for OLMo-1B, leveraging Fully Sharded Data Parallel (FSDP) for continual pre-training. For LLaMA, we adopted the LLaMA-Factory framework to carry out the continual pre-training process. Our experiments was primarily conducted on OLMo-1B and Llama-3-8B models, with Llama-3-8B utilizing LoRA (Low-Rank Adaptation) for parameter-efficient fine-tuning. The data and evaluation are given in this repo [^3] https://github.com/microsoft/LMOps/tree/main/adaptllm. We conducted the continual pre-training on a total of 1B tokens.
@@
We used the Llama-Factory (Zheng et al., 2024) framework to fine-tune Llama-3-8B. As for general instruction tuning tasks, we adopt instruction tuning datasets from (Xia et al., 2024)  [^4] https://huggingface.co/datasets/princeton-nlp/less_data, including CoT (Wei et al., 2022) , FLAN v2 (Longpre et al., 2023), and Open Assistant 1 (Kopf et al., 2023). As for code-related reasoning tasks, we utilize OSS-Instruct-75K  [^5] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K and Evol-Instruct-110K  [^6] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K. These datasets provide sufficient diversity for verification on fine-tuning. We apply LoRA (Hu et al., 2021) to Llama-3-8B experiments.
F.2 Evaluation
@@
Table 13: PPL results of GPT-2 124M pre-training on mixture of human and synthetic data.
@@
Appendix G Discussion
G.1 Non-Iterative vs Iterative Model Collapse.
The difference between the two lies in their scope. Non-iterative model collapse is not confined to training on self-generated data, allowing it to uncover broader properties of synthetic data. For instance, in our experiments, we train GPT-2 on the Cosmopedia dataset in a single generation, which was generated by Mixtral-8x7B-Instruct-v0.1. In contrast, iterative model collapse focuses on training the model over multiple generations using self-generated data.
Furthermore, the non-iterative model collapse emphasizes the gap between human data and general purely synthetic data, particularly regarding distributional properties and n-gram features. In contrast, the iterative model collapse illustrates the iterative evolution of the model, resembling a self-play process. This process illustrates the gradual evolution of self-generated data. It does not involve an analysis of the differences in nature between self-generated and human-produced data. They both ultimately lead to model collapse, driven by the same underlying cause—synthetic data, though they investigate different aspects of synthetic data. We often face that training a model on a mixture of human and synthetic data, where the synthetic data is not generated by the model itself, and its exact origin may be unknown.
G.2 Why Does the Observed Probability Distribution Exhibit Filtering Potential?
From the perspective of information theory, we can analyze the filtering potential of the U-shape distribution as follows: We utilize the U-shape distribution in Figure 5 to re-sample tokens in the high-probability region, to adjust the U-shaped distribution toward a uniform distribution. By doing so, we can maximize the information entropy. According to information theory, maximizing information entropy is achieved when the distribution is uniform.
Lemma 1: Let $X$ be a discrete random variable with $n$ possible outcomes. If the probability of each outcome is uniform, i.e., $P(x_{i})=\frac{1}{n}$ for all $i\in\{1,2,\dots,n\}$, the Shannon entropy is maximized, given by:
$H(X)$ $=-\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n.$ (45)
G.3 Gradual Decline in Editing
Table 15: Percentage of tokens requiring edits in the Natural-Instructions dataset. The total number of tokens is 4,671,834.
Gen 1 (source) Gen 2 Gen 3
Tokens ($p>0.99$) 584,103 549,519 517,433
Percentage 12.5% 11.76% 11.08%
We present the percentage statistics of edited tokens in Table 15 and performance in an iterative process in Table 16, demonstrating that the edited tokens indeed exhibit a progressive decrease. Specifically, We observe that the percentage of edited tokens (above the threshold $p>0.99$) decreases as the generation number increases. Theoretically, this is a process of distribution shifting. When tokens ($p>0.99$) are resampled, randomness is introduced. The sampling process can select tokens with lower probabilities. Then, tokens ($p>0.99$) is replaced, leading to a reduction of edited tokens in subsequent generations. The Table 15 provides empirical evidence for this pattern of decay.
Table 16: Performance in an iterative process on Instruction tuning data.
PIQA BoolQ HS SIQA WG Avg
Gen 0 79.87 81.28 59.72 49.69 74.51 69.01
Gen 1 80.25 81.16 59.74 50.56 74.59 69.26
Gen 2 80.14 82.69 59.82 50.51 73.80 69.39
G.4 What is Coverage Narrowing?
‘coverage narrowing’ refers to a phenomenon in which the distribution of synthetic data covers a significantly narrower range of values compared to human data, even when the data sizes are identical. For instance, as shown in Figure 3, the PPL range of synthetic data is limited to $[0,14]$, whereas the PPL range of human data extends from $[0,100+]$. Despite this disparity, the overall coverage, represented by the area under the distribution curves, remains the same. This significant distribution gap is what we define as ‘coverage narrowing.’
G.5 How Does the DSIR Work?
@@
G.6 Non-autoregressive Token Replacement May Compromise Text Coherence.
@@
G.7 Must We Assume the Data is 100% Human-authored?
@@
G.8 Can ToEdit Help with Already Strongly Collapsed Data or is a Minimum Quality of the Data Necessary?
The ToEdit algorithm was initially designed to preserve the long-tail distribution during the data generation process, thereby avoiding model collapse. For already collapsed data, the variance is typically very small, and enhancing diversity is crucial. We can also adjust the threshold to introduce more randomness in data. Through this operation, we can inject randomness into the collapsed data. However, this is a theoretical scenario, and as we know, data situations are highly complex. In practice, there will be many more challenges to address.
G.9 Could We Generate Tons of Data Using LLMs and Select Only the Long-Tail Ones?
For now, generating long-tail samples is currently difficult for language models. The reason lies in the sampling strategy of LLMs. Current LLMs adopt top-p, top-k or other sampling strategy for better performance. However, these sampling strategy will lead to cut-off distribution. When the data synthesizing scale up, this drawback will finally scaling law cut-off on synthetic data. However, human corpus data follows a Zipf distribution. The truncated output distribution causes the LLMs to nearly fail to sample long-tail samples. In other words, it is currently difficult to induce long-tail samples from LLMs that are as diverse as human data.
On the other hand, if we force the language model to generate long-tail samples, these may contain both noisy and high-information samples, which are like two sides of a coin, both distributed in the long tail of the data. This necessitates further filtering of the high-information samples. Unfortunately, such samples are challenging to automatically identify in practice and may require extensive human annotation.
Appendix H Potential Applications and Future Work
@@
Extended related applications
A broader range of recent works across domains also explore synthetic data generation and usage for diverse applications (Bai et al., 2025, 2024, 2023; Jia et al., 2024; Yang et al., 2023; Li et al., 2020; Niu et al., 2024; Zhang et al., 2023; Li et al., 2024; Pu et al., 2024; Li et al., 2023; xu2020siamfc; Zhu et al., 2024a, b, 2022).